OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4). y = four thirdsx + 8 y = four thirdsx + 3 y = -three fourthsx + 8 y = -three fourthsx + 3 @AriPotta

OpenStudy (aripotta):

ok first convert 4x - 3y = -12 into slope-intercept form

OpenStudy (anonymous):

Slope intercept form is \[y=mx+b\]

OpenStudy (anonymous):

I don't know how :(

OpenStudy (aripotta):

ok so putting it in slope-intercept form is basically solving for y. so we want y by itself. first, we're going to subtract 4x from both sides -3y = -4x - 12 then divide both sides by -3 to get y = 4/3x + 4

OpenStudy (aripotta):

parallel lines have the same slope

OpenStudy (aripotta):

so since we know the slope of the line, and a point the line goes through, (-3,4), we can use the point-slope form: y - y1 = m(x - x1)

OpenStudy (aripotta):

so we're going to plug the info into the formula: y - 4 = 4/3(x + 3)

OpenStudy (aripotta):

then we have to convert it into slope-intercept form y - 4 = 4/3(x + 3) y - 4 = 4/3x + 4 y = 4/3x + 8

OpenStudy (aripotta):

make sense?

OpenStudy (anonymous):

kinda yeah

OpenStudy (aripotta):

ok cool :)

OpenStudy (aripotta):

do you have any questions?

OpenStudy (anonymous):

thats it, thank you

OpenStudy (aripotta):

\(\Large \boldsymbol{ \rlap{\color{#008AFF}{\hspace{2.5pt}n}\color{#0073FF}{\hspace{0.34pt}o} \hspace{3pt}\color{#005CFF}{p} \color{#0045FF}{\hspace{0.9pt}r} \color{#0017FF}{o} \color{#0200FF}{\hspace{1.7pt}b} \color{#3000FF}{\hspace{1.5pt}l} \color{#4700FF}{\hspace{0.4pt}e} \color{#7500FF}{\hspace{0.9pt}m} \color{#8C00FF}{\hspace{0.5pt}o}~\LARGE\color{#D100FF}{\ddot\smile}}{\hspace{1.5pt} no~p\hspace{0.3pt}r\hspace{0.4pt}o\hspace{1pt}b\hspace{1.5pt}l\hspace{0.4pt}e\hspace{0.4pt}m\hspace{0.4pt}o}}\)